OpenStudy (anonymous):

In music, a dotted note has $1\dfrac{1}{2}$ the duration of the original note. What fraction of a whole note is a dotted eighth note? Note: An eighth note is one-eighth of a whole note (as the name suggests).

OpenStudy (ibbutibbu.):

What is $1\dfrac{1}{2}$ ?

OpenStudy (anonymous):

wow this is advanced sorry i cant help with it

OpenStudy (anonymous):

it actually isn't as hard it's supposed to be 1 and 1/2 the $1\dfrac{1}{2}$

OpenStudy (anonymous):

$1\dfrac{1}{2}$=1 and 1/2

OpenStudy (anonymous):

do you get it?